Which function describes the data shown below in the table?
x: -6, 0, 6, 12
y: 81, 27, 9, 3
Is it (a) y=81(3)^x/6 (b) y=81(1/3)^x/6 (c) y= 27(1/3)^6x (d) y= 27(1/3)^x/6? I think it is (b) ????????

Answer by ilana(307)

Nope, not (b). You could test that answer by plugging in any of the x values and seeing if you get the right y. I use 0 because I like when life is easy. (b) is y=81(1/3)^x/6. So if x=0, y=81(1/3)^0/6 = 81(1)=81. But the table says that when x=0, y=27. So that didn't work. The answer is (d). You got the rate (1/3) and the exponent (x/6) right. You just got tricked by the fact that the table starts at x=-6 instead of at x=0.